t=5(logP/log0.5)
P = percent
t = 9:00a.m.
Therefore equation is:
9=5(logP/log0.5)
What is the percent?

Answer by Theo(13342)

start with 9 = 5 * (log(p) / log(.5)
divide both sides of the equation by 5 to get:
9/5 = log(p) / log(.5)
multiply both sides of the equation by log(.5) to get:
9/5 * log(.5) = log(p)
solve for log(p) to get:
log(p) = -.5418539922
this is true if and only if 10^(-.5418539922) = p
solve for p to get:
p = .2871745887